OpenStudy (anonymous):

x varies directly with y and z. x = 1200 when y = 20 and z = 30. Find x when y = 10 and z = 20.

OpenStudy (luigi0210):

\[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (anonymous):

can you help me out

OpenStudy (luigi0210):

If I'm not mistaken, the equation for this is: x=kyz

OpenStudy (anonymous):

yes ur correct

OpenStudy (luigi0210):

Okay, so plug in the first set of values

OpenStudy (anonymous):

how is the so is it 1200=k(20)(30)

OpenStudy (luigi0210):

Now solve for k

OpenStudy (anonymous):

is 20

OpenStudy (luigi0210):

20 or 2?

OpenStudy (anonymous):

2

OpenStudy (luigi0210):

Good, now that you have k, plug in the second set of values and you are done good sir :)

OpenStudy (anonymous):

x=2(10)(20)

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

400

OpenStudy (luigi0210):

Congrats, you have found your answer.

OpenStudy (luigi0210):

And yes, it is that easy :)